Circular No. 8030

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002jm IN IC 603 Further to IAUC 8026, M. Ganeshalingam and W. Li report the dis-

covery by LOTOSS of an apparent supernova (mag ∼ 17.5) on an unfiltered KAIT image taken on Dec. 7.5 UT, confirmed at about the same brightness by M. Schwartz with the 0.8-m Tenagra II telescope on Dec. 9.5. SN 2002jm is located at α = 10^{h}19^{m}25^{s}.20, δ = −5^{o}39′14′′.5 (equinox 2000.0), which is 2′′.6 east and 7′′.3 north of the center of IC 603. A KAIT image taken on Nov. 23.5 showed nothing at this position (limiting mag ∼ 18.5).

SUPERNOVA 2002jn IN UGC 11523 Further to IAUC 8009, T. Boles reports the discovery of an apparent

supernova (mag 18.0) on an unfiltered CCD image taken on Dec. 9.767 UT with a 0.35-m reflector. The new object is located at α = 20^{h}11^{m}59^{s}.91, δ = +5^{o}31′16′′.0, which is ≈ 10′′.0 west and 11′′.5 north of the center of UGC 11523. M. Schwartz, Cottage Grove, OR, confirmed the new object with the 0.82-m Tenagra telescope on Dec. 11.073. Boles measured position end figures 59^{s}.94, 15′′.6 from Schwartz’ image. SN 2002jn is not present on Boles’ images from Oct. 18 (limiting mag 18.5) or Sept. 28 (limiting mag 19.0), and it is not present on Palomar Sky Survey red and blue plates from 1990 or on a Quick-V survey plate from 1982.

COMET C/2002 X1 (LINEAR) Additional astrometry, now spanning Dec. 5–9, produces the following

revised parabolic orbital elements (MPEC 2002-X58):

T = 2003 July 12.733 TT  ω = 206.^{o}938
Ω = 281.855 }2000.0
q = 2.49886 AU  i = 164.111
2002/03 α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Dec. 2 10^{h}36.^{m}81 +19^{o}28.^{′}5 3.183 3.429 96.^{o}0 16.^{o}6 15.9
12 10 27.73 +20 32.6 2.919 3.361 108.4 16.1 15.6
22 10 14.25 +21 57.4 2.669 3.295 121.8 14.7 15.3
Jan. 1 9 55.30 +23 41.1 2.447 3.230 136.3 12.1 15.0
11 9 30.04 +25 35.6 2.265 3.166 151.9  8.4 14.8
21 8 58.50 +27 24.3 2.138 3.104 166.8  4.1 14.6
31 8 22.35 +28 45.1 2.077 3.044 166.9  4.2 14.4
2002 December 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green